40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jun 19 18:56:55 2018, Jon, Update, General, AUX beam alignment issues 
    Reply  Wed Jun 20 00:57:04 2018, johannes, Update, General, AUX beam alignment issues as110_rf_30MHz.pdfyend_rf_30MHz.pdf
       Reply  Sun Jun 24 23:13:46 2018, johannes, Update, General, AUX beam alignment issues AS_AUX_SETUP.pdf
Message ID: 13987     Entry time: Tue Jun 19 18:56:55 2018     Reply to this: 13989
Author: Jon 
Type: Update 
Category: General 
Subject: AUX beam alignment issues 

Not much progress today with the AUX cavity scans. I've determined there still are some alignment issues.

At the start of today a large AUX/PSL beat note was visible on the AS110 sensor, at a similar power as where we left off last night (-60 dBm). Proceeding from there, I attempted to reproduce Johannes' measurement of the cavity transmission resonances. I misaligned the X-arm, locked the Y-arm cavity, and scanned the AUX RF offset approximately 8 MHz in 2 kHz steps. This should have swept through two FSRs, but nothing was visible.

Further inspection revealed that none of the PSL light was making it back to through the AUX fiber to the PSL table. I take this to mean that the beam seen earlier on AS110 was the ITMY reflection, and that the AUX injection axis was no longer reaching ETMY. I also found that the AUX beam size just after the 90/10 beasmsplitter looks anomolously large. Maybe a lens was recently changed? In any case, the mode-matching looks like it is going to need to be readjusted.

ELOG V3.1.3-